OpenStudy (anonymous):

PLEASE can somebody check my work? Only 5 questions. (involving sohcahtoa) I really need to ace this. -.-

OpenStudy (anonymous):

Question 1 (Fill-In-The-Blank Worth 4 points) (07.02 MC) Use △DEF, shown below, to answer the question that follows. Triangle DEF where angle E is a right angle. DE measures x. DF measures 55. Angle D measures 49 degrees. What is the value of x rounded to the nearest hundredth? Type the numeric answer only in the box below. Answer for Blank 1: Question 2 (Multiple Choice Worth 4 points) (07.02 MC) Given the triangle below, which of the following is a correct statement? Right triangle ABC with AB measuring 5, AC measuring 6, and BC measuring 8. cot right angle triangle B equals fraction 6 over 5 csc right angle triangle C equals fraction 3 over 4 cot right angle triangle C equals fraction 8 over 5 csc right angle triangle b equals fraction 4 over 3 Question 3 (Multiple Choice Worth 4 points) (07.02 MC) In △PQR, find the measure of ∡P. Triangle PQR where angle Q is a right angle. PQ measures 33 point 8. PR measures 57 point 6. Measure of angle P is unknown 30.4° 59.6° 35.9° 54.1° Question 4 (Multiple Choice Worth 4 points) (07.02 MC) In △JKL, solve for x. Triangle JKL where angle K is a right angle. KL measures 34. JK measures x. Angle J measures 27 degrees. 17.32 38.16 66.73 74.89 Question 5 (Multiple Choice Worth 4 points) (07.02 MC) Find the value of the sec 19° using your calculator. 0.946 1.011 0.989 1.058

OpenStudy (anonymous):

Answers: 36.08 D D A D

OpenStudy (anonymous):

Ignore where it says (7.02 mc)

OpenStudy (anonymous):

number one^

OpenStudy (anonymous):

number 2

OpenStudy (anonymous):

3

OpenStudy (anonymous):

4

OpenStudy (lastdaywork):

Q2 - I find all incorrect Q4 - NOT A

OpenStudy (lastdaywork):

Given the triangle below, which of the following is a correct statement? Right triangle ABC with AB measuring 5, AC measuring 6, and BC measuring 8. cot right angle triangle B equals fraction 6 over 5 csc right angle triangle C equals fraction 3 over 4 cot right angle triangle C equals fraction 8 over 5 csc right angle triangle b equals fraction 4 over 3 @texaschic101 Please answer this ^^ question. I found all options to be wrong.

OpenStudy (phi):

your answers look good except for Q4,

OpenStudy (anonymous):

ok, thank you! I re-did number 4 and got answer C. Is that right?

OpenStudy (anonymous):

C is right, for anybody who finds this article later. :)

OpenStudy (phi):

yes,4 C

OpenStudy (lastdaywork):

@phi Please explain me Q2 ??

OpenStudy (lastdaywork):

*Omit ?? XD

OpenStudy (phi):

Right triangle ABC with AB measuring 5, AC measuring 6, and BC measuring 8. cot right angle triangle B equals fraction 6 over 5 csc right angle triangle C equals fraction 3 over 4 cot right angle triangle C equals fraction 8 over 5 csc right angle triangle b equals fraction 4 over 3 First, notice they cheated. They say the longest side (hypotenuse) is 8 in a right triangle we know a^2 + b^2 = c^2 so we expect 5^2 + 6^2 = 8^2 --> 25+36= 64 or 61=64 which is close, but not true.